OpenStudy (anonymous):

I need help finding the value "a" from 3a + 2=a - 6

OpenStudy (chaise):

3a + 2=a - 6 Take away two from both sides of the equal signs 3a = a -8 Take away a from both sides of the equal signs 2a = -8 Divide by two from both sides of the equal sign a = -4

OpenStudy (anonymous):

first subtract a from both sides of the equation: 3a - a + 2 = a - a - 6 2a + 2 = -6 subtract 2 from both sides 2a = - 8 divide both sides by 2: a = -4